What mass of KX (Mm= 190 g/mol) is required to prepare 555 mL of a pH = 12.10 solution? The pKa of HX is 12.82

___ g KX
Answer #1

HX +420 ② x +H₂O (x=kx ] x tho HX + oH giren pH=12-10 pt + polt=14 PO 14-12.10 = 1.9 [olf] = 10la = 0012619 at Equilibrium (- moles molanly = noen X1ooo Volume molarity = weight x 100 M.Wfx volume weight = 11 molantas XM. wt x volume 1000 = 0.363 mo